OpenStudy (anonymous):

Joey has a collection of baseball cards that he has separated into two different value categories. His common cards are worth $0.75 each and his premium cards are worth $2.00 each. Together, Joey’s entire baseball card collection of 88 cards is worth $106.00. How many cards in Joey’s collection are worth $0.75?

OpenStudy (texaschic101):

c + p = 88 --> p = 88 - c 0.75 + 2 = 106 now sub 88 - c in for p in the 2nd equation and solve for c (common cards worth 75 cents) and you will have your answer :)